# Radiant Triangle Soup with Soft Connectivity Forces for 3D Reconstruction and Novel View Synthesis

### v1

@@ 4.3 Diffuse Primitive Boundaries
Alpha-blending is the primary driving force of optimization for all splatting frameworks. Blending across primitives facilitates optimization via gradient descent and encourages, among other behaviors, movement in primitive position and orientation. It is challenging in practice to optimize a scene with fully opaque primitives that do not smoothly blend with one another. It is therefore ideal to make the triangle primitives diffuse near the edges. Taking inspiration from 3DCS [5], we parameterize the response from each triangle as a function of the signed distance from the ray-triangle intersection to the nearest edge,
@@
where $w_{\sigma}$ is the diffuse weight, $l$ is the signed distance between the intersection and the nearest triangle edge in the plane of the triangle, and the scalar $\sigma$ is an optimizable parameter of each triangle that controls the level of diffuseness. As $\sigma$ increases, the boundary of the triangle becomes less diffuse, creating sharper primitive renderings (see Fig. 4). For the signed distance, $l<0$ occurs when the intersection point lies outside of the triangle boundary.
@@ 4.4 Rasterization
With the barycentric coordinates of the ray-triangles intersections and diffuse weight computed, we interpolate the color for the current pixel w.r.t. the colors of each vertex weighted by the barycentric coordinates,
$c_{n}=\begin{bmatrix}c^{0}&c^{1}&c^{2}\end{bmatrix}\lambda$ (6)
where $c^{0}$, $c^{1}$, and $c^{2}\in\mathbb{R}^{3}$ are computed from the SH components of each vertex, and $\lambda$ are the barycentric coordinates of the intersection point.
We aggregate the contribution to the current output pixel, $i$, from each intersected primitive,
$c_{i}=\sum w_{n}c_{n}$ (7)
where $w_{n}=\alpha w_{\sigma}T$, and $T=\Pi_{j=1}^{i-1}(1-\alpha_{j})$ is the transmittance for the current primitive [1].
Previous works [2, 4, 13] provide two methods for rendering per-pixel depth; (1) computing the average weighted intersection depth of all traversed primitives (mean depth), and (2) using only the depth of the primitive that causes the transmittance $T$ to exceed $0.5$ (median depth). In our work, we use median depth. Using the average weighted depth encourages the formation of many semi-transparent layers of primitives. Rendering median depth removes this blending and helps guide the formation of surfaces. We directly use $d$ from Eq. 2 for the depth of each intersection.
For each triangle, we compute the surface normal as the cross-product between two edges,
$\widehat{n}_{n}=\frac{\overrightarrow{V^{0}V^{2}}\times\overrightarrow{V^{0}V^{1}}}{||\overrightarrow{V^{0}V^{2}}\times\overrightarrow{V^{0}V^{1}}||_{2}}$ (8)
Unlike the rendered depth, to render per-pixel surface normals, we follow previous work [2, 4] and alpha-blend all the surface normals of all intersected primitive. The intuition is that through a normal consistency loss (see Section 5), the blended normals must align with the normals computed from the median depth map. This encourages all the intersected triangles to align (and eventually collapse onto) the median surface. Thus, we render per-pixel surface normals through alpha-blending,
$\overrightarrow{n}_{i}=\sum w_{n}\widehat{n}_{n}$ (9)
@@ 4.5 Connectivity
To enable soft connectivity among primitives and encourage flat, connected surfaces, we add connectivity and surface normal terms to the optimization objective. For each triangle edge, we assign, and periodically update, a connection to the nearest neighboring triangle edge. Taking inspiration from the energy functions introduced in 3D scene flow estimation [28], the loss of the connectivity increases according to the distance between their vertices and the inner product between their normals, discussed in Section 5. In order to prevent primitives from undesired connections, edge connection assignment must pass direction and orientation criteria, shown in Fig. 5. Simply, we must ensure that the neighboring triangle edge is in the correct direction and is oriented appropriately in order to connect with the current edge. We compute connectivity assignment and activate the connectivity losses after several epochs.
Refer to caption Refer to caption
(a) Direction Constraint (b) Orientation Constraint
Refer to caption Refer to caption
(c) $\tau$ (d) $\rho$
Figure 5: Above we show connectivity constraints between two triangles. For simplicity, it is assumed the triangles are co-planar.
Naively connecting with the nearest edge without any consideration of relative position may lead to connections that cross triangles and could cause them to flip. To prevent such connections, we compute three unit vectors: (1) the unit vector orthogonal to the current triangle edge and the current triangle normal (pointing outside the triangle), (2) the unit vector orthogonal to the neighboring triangle edge and the neighboring triangle normal (pointing outside the triangle), and (3) the connecting vector from the center of the current triangle edge to the center of the neighboring triangle edge (normalized to be unit length). To measure the relative position between the triangles, we take the inner product between these unit vectors. The direction criterion is measured as the inner product between the current triangle edge vector and the connecting vector, shown in Fig. 5(c). The orientation criterion is measured as the inner product between the current triangle edge vector and the neighboring triangle edge vector, shown in Fig. 5(d). We apply thresholds, $\rho$ and $\tau$ as shown in the figure, to these two values to prevent any unwanted connectivity from forming.
Intuitively, the direction criterion ensures the neighboring edge is in the correct direction w.r.t. the current triangle. Not satisfying this criterion would cause the current triangle to flip in order to reduce the connectivity loss. The orientation criterion ensures the edges are facing each other and the triangles have relatively similar surface normals. Not upholding this criterion would cause the neighboring triangle to flip in order to reduce the connectivity loss.
4.6 Adaptive Densification
Figure 6: During densification, large triangles are split into four sub-triangles with interpolated colors per-vertex.
@@ 4.6 Adaptive Densification
We perform adaptive density control through the process of cloning and splitting triangles. Previous works [2, 1, 4, 3] perform the cloning and splitting procedures by duplicating primitives conditioned on scale and position gradients. In order to properly split large triangles, we must split them into four sub-triangles. During early iterations in the optimization, when triangles are split, the sub-triangles move independently to better align with surfaces in the scene. During the later stages when most of the surfaces have formed, the triangles split to enable more detailed surface color representation, remaining relatively attached via the connectivity forces. See Fig. 6 for a visualization of the splitting procedure. Similar to previous work, we directly clone small primitives selected for densification.
@@
Much like 2DGS [2], our method does not directly rely on the gradient of the projected 2D primitive center. Instead of computing an approximation via projecting 3D gradients into the camera plane [2], we directly condition densification on the magnitude of the reference point gradient $\nabla\mu$ in 3D.
@@ 5 Optimization
Our objective function comprises terms computed in 2D, on the images planes, and in 3D, on the triangles.
5.1 Rendering Losses
Following previous work [1, 2, 3, 4, 13], we compute the SSIM between the rendered and input images,
$\mathcal{L}_{ssim}=\frac{1}{N}\sum_{i,j}(1-\gamma)L_{1}+\gamma L_{ssim}$ (10)
To help locally align the triangles with the rendered surface, we apply a normal consistency loss [2]. This encourages alignment between the rendered normal $\widehat{n}$ with the normal computed from the rendered depth map $\widehat{n}_{d}$,
$\mathcal{L}_{norm}=\frac{1}{N}\sum_{i,j}(1-\widehat{n}^{T}\widehat{n}_{d})$ (11)
From the unsupervised depth estimation literature [29, 30, 31, 32, 33], we adopt a smoothness term on the rendered depth map conditioned on the gradient of the input image. This penalizes large gradients in the rendered depth map where we have small gradients in the input image,
$\mathcal{L}_{smooth}=\frac{1}{N}\sum_{i,j}||\partial_{x}D_{i,j}||e^{||\partial_{x}I_{i,j}||}+||\partial_{y}D_{i,j}||e^{||\partial_{y}I_{i,j}||}$ (12)
5.2 Scene Loss
$\mathcal{L}_{conn}=\sum_{a\in\Omega}\frac{1}{2}(||V_{a}^{1}-V_{b}^{1}||_{2}+||V_{a}^{2}-V_{b}^{2}||_{2})+(1-\widehat{n}_{a}^{T}\widehat{n}_{b})$ (13)
where $V_{a}^{j}$ and $V_{b}^{j}$ are the $j^{th}$ vertex pair of connected edges $a$ and $b$, respectively. $\widehat{n}_{a}$ and $\widehat{n}_{b}$ are the normals of the connecting triangles and $\Omega$ is the set of all triangles that intersect the current camera frustum. Activating this loss for invisible triangles without rendering losses leads to oversmoothing.
The final objective is a weighted summation of the above,
6 Experiments
Table 1: Chamfer distance evaluation on scenes from DTU [6]. Following previous literature, we average the accuracy and completeness Chamfer distances on the widely used evaluation set. Chamfer distances are measured in millimeters.
@@ 6 Experiments
24 37 40 55 63 65 69 83 97 105 106 110 114 118 122 Mean (mm)$\downarrow$
3DGS [1] 2.14 1.53 2.08 1.68 3.49 2.21 1.43 2.07 2.22 1.75 1.79 2.55 1.53 1.52 1.50 1.96
SuGaR [3] 1.47 1.33 1.13 0.61 2.25 1.71 1.15 1.63 1.62 1.07 0.79 2.45 0.98 0.88 0.79 1.33
2DGS [2] 0.48 0.91 0.39 0.39 1.01 0.83 0.81 1.36 1.27 0.76 0.70 1.40 0.40 0.76 0.52 0.80
GOF [13] 0.50 0.82 0.37 0.37 1.12 0.74 0.73 1.18 1.29 0.68 0.77 0.90 0.42 0.66 0.49 0.74
PGSR [4] 0.36 0.57 0.38 0.33 0.78 0.58 0.50 1.08 0.63 0.59 0.46 0.54 0.30 0.38 0.34 0.52
RTS 0.48 0.59 0.79 0.39 0.58 0.92 0.77 0.94 0.77 0.37 0.79 1.12 0.37 0.71 0.49 0.67
6.1 Implementation Details
We implement the majority of our RTS framework in Python using PyTorch [34]. For rasterization, we develop custom CUDA kernels for both the forward and backward pass. We run all our experiments on a single NVIDIA RTX A6000. In all of our experiments, we choose $\tau=0.0$ and $\rho=0.0$ for the threshold values on the connectivity constraints. For our loss weights, we choose $\omega=\begin{bmatrix}1.0&0.05&0.8&10.0\end{bmatrix}$ empirically.
Following state-of-the-art Multi-View Stereo methods [35, 36, 37, 38, 39, 40], we directly generate a final 3D point cloud from the rendered depth maps. To generate our final point cloud, we use heuristic filtering, similar to the post-processing presented in GBiNet [40]. For each rendered depth map, we project each depth estimate into neighboring views, using the depth estimates in each view at the projected pixel location to reproject back to the reference view, measuring the pixel reprojection error and filtering out estimates whose error is above a threshold.
6.2 Evaluation
We test our framework on DTU [6], which is an indoor dataset that contains images of 124 scenes taken from a camera mounted on an industrial robot arm. All scenes share the same camera trajectories, with ground-truth point clouds captured via structured light.
We evaluate our new approach on the DTU dataset and record the Chamfer distance in Table 1. We show competitive results alongside the leading state-of-the-art SGS methods. Across all scenes in the test set, we report highly competitive metrics against the current state of the art. Notably, our method seems particularly sensitive to specific scenes with texture-less and non-Lambertian surfaces. Qualitative results can be seen in Fig. 7.
@@ 6.2 Evaluation
Figure 7: Qualitative results on several scenes from the DTU dataset [6].
We measure the quality of novel view synthesis for our method on the DTU dataset using standard metrics of PSNR, SSIM, and LPIPS [41] in Table 2. Our method produces comparable results, with a slight reduction in PSNR. As von Lützow and Nießner [25] note, using primitives with explicit boundaries can begin to introduce hard edges in regions with poor visibility. While the diffuse boundaries of our triangles helps prevent much of this behavior, its effects are noticeable in some of the reconstructed images.
Table 2: Photometric evaluation on several scenes from the DTU [6] dataset. We record the per-scene SSIM, PSNR, and LPIPS metrics on the widely used test set.
Method DTU
@@
SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$
SuGaR [3] - 34.57 -
2DGS [2] 0.939 34.33 0.169
PGSR [4] 0.917 33.18 0.251
RTS 0.918 31.68 0.213

### v2

@@ 3.3 Diffuse Primitive Boundaries
Alpha-blending is the primary driving force of optimization for all splatting frameworks. Blending across primitives facilitates optimization via gradient descent and encourages, among other behaviors, movement in primitive position and orientation. It is challenging in practice to optimize a scene with fully opaque primitives that do not smoothly blend with one another. It is therefore desirable to make the triangle primitives diffuse near the edges. Taking inspiration from 3DCS (Held et al., 2025b), we parameterize the response from each triangle as a function of the signed distance from the ray-triangle intersection to the nearest edge,
@@
where $w_{\sigma}$ is the diffuse weight, $l$ is the signed distance between the intersection and the nearest triangle edge in the plane of the triangle, and the scalar $\sigma$ is an optimizable parameter of each triangle that controls the level of diffuseness. As $\sigma$ increases, the boundary of the triangle becomes less diffuse, creating sharper primitive renderings (see Fig. 4). For the signed distance, $l<0$ occurs when the intersection point lies outside of the triangle boundary. This formulation slightly deviates from that of 3DCS, as the signed distance is computed directly, compared to their smooth approximation.
@@ 3.4 Rasterization
Throughout this paper, we use barycentric coordinates, $\lambda=[\lambda^{0},\lambda^{1},\lambda^{2}]^{T}$, to rasterize triangles. Please see Section S.1 for more details. With the barycentric coordinates of the ray-triangle intersections and the diffuse weight computed, we interpolate the color for the current pixel w.r.t. the colors of each vertex weighted by the barycentric coordinates, $c_{n}=\begin{bmatrix}c^{0}&c^{1}&c^{2}\end{bmatrix}\lambda$, where $c^{0}$, $c^{1}$, and $c^{2}\in\mathbb{R}^{3}$ are computed from the SH components of each vertex, and $\lambda$ are the barycentric coordinates of the intersection point.
We aggregate the contribution to the current output pixel, $i$, from each intersected primitive, $c_{i}=\sum w_{n}c_{n}$, where $w_{n}=\alpha w_{\sigma}T$, and $T=\Pi_{j=1}^{i-1}(1-\alpha_{j})$ is the transmittance for the current primitive (Kerbl et al., 2023).
Previous works (Huang et al., 2024; Chen et al., 2024a; Yu et al., 2024) provide two methods for rendering per-pixel depth; (1) computing the average weighted intersection depth of all traversed primitives (mean depth), and (2) using only the depth of the primitive that causes the transmittance $T$ to exceed $0.5$ (median depth). In our work, we use median depth. Using the mean depth encourages the formation of many translucent layers of primitives. Rendering median depth removes this blending and helps guide the formation of surfaces. We directly use $d$ from Eq. S.2 for the depth of each intersection.
Additionally, we compute the surface normal, $\widehat{n}_{n}$, for each triangle as the cross-product between two edges. Unlike the rendered depth, to render per-pixel surface normals, we follow previous work (Huang et al., 2024; Chen et al., 2024a) and alpha-blend all the surface normals of all intersected primitives. The intuition is that through a normal consistency loss (see Section 4), the blended normals must align with the normals computed from the median depth map. This encourages all the intersected triangles to align (and eventually collapse onto) the median surface. Thus, we render per-pixel surface normals through alpha-blending, $\overrightarrow{n_{i}}=\sum w_{n}\widehat{n}_{n}$.
Figure 5: Examples of edge association for connecting nearby triangles. For simplicity, it is assumed the triangles are co-planar in this figure.
@@ 3.5 Connectivity
To enable soft connectivity among primitives and encourage flat, connected surfaces, we add a connectivity term to the optimization objective. For each triangle edge, we assign, and periodically update, a connection to the nearest neighboring triangle edge. Taking inspiration from the energy functions introduced in 3D scene flow estimation (Vogel et al., 2015), the connectivity term of the loss increases according to the distance between their vertices and the inner product between their normals, discussed in Section 4.
Naively connecting with the nearest edge without considering the relative orientations may cause connections that would require large changes in rotation to either triangle, leading to undesirable behaviors during optimization. To prevent this behavior, connections are only established if the triangle edges are "facing" each other. In practice, we use the inner product between the unit vectors orthogonal to the triangle edges (in the plane of each triangle) as the criterion for establishing connections. We provide an example in Fig. 5 where a connection to the highlighted edge of triangle D is valid, while the other two connections would cause large rotations.
3.6 Adaptive Density Control
@@
We perform adaptive density control through the process of cloning, splitting and pruning triangles. Previous works (Kerbl et al., 2023; Huang et al., 2024; Guédon & Lepetit, 2024; Chen et al., 2024a) perform the cloning and splitting procedures by duplicating primitives conditioned on scale and position gradients. In order to properly split large triangles, we must split them into four sub-triangles. During early iterations in the optimization, when triangles are split, the sub-triangles move independently to better align with surfaces in the scene. During the later stages when most of the surfaces have formed, the triangles split to enable more detailed surface color representation, remaining relatively attached via the connectivity forces. See Fig. 6 for a visualization of the splitting procedure.
Figure 6: Triangles are split with interpolated colors per-vertex.
@@
Similar to previous work, we directly clone small primitives selected for densification. Much like 2DGS (Huang et al., 2024), our method does not directly rely on the gradient of the projected 2D primitive center. Instead of computing an approximation via projecting 3D gradients into the camera plane (Huang et al., 2024), we directly condition densification on the magnitude of the incenter gradient $\nabla\mu$ in 3D.
To remove uninformative triangles, we prune primitives that meet the following criteria: (i) triangles that are transparent ($\alpha<0.05$), (ii) triangles with one or fewer edge connections, (iii) triangles that do not intersect the camera frustum of at least three views after each epoch.
@@ 4 Optimization
Our objective function comprises terms computed in 2D (on the images plane), and in 3D (on the triangles).
Table 1: Chamfer distance evaluation on scenes from DTU (Aanæs et al., 2016). Following previous literature, we average the accuracy and completeness Chamfer distances on the widely used evaluation set. Chamfer distances are measured in millimeters. The best results are in boldface and the second best are underlined.
@@
[…]
3DGS (Kerbl et al., 2023) 2.14 1.53 2.08 1.68 3.49 2.21 1.43 2.07 2.22 1.75 1.79 2.55 1.53 1.52 1.50 1.96
SuGaR (Guédon & Lepetit, 2024) 1.47 1.33 1.13 0.61 2.25 1.71 1.15 1.63 1.62 1.07 0.79 2.45 0.98 0.88 0.79 1.33
2DGS (Huang et al., 2024) 0.48 0.91 0.39 0.39 1.01 0.83 0.81 1.36 1.27 0.76 0.70 1.40 0.40 0.76 0.52 0.80
GOF (Yu et al., 2024) 0.50 0.82 0.37 0.37 1.12 0.74 0.73 1.18 1.29 0.68 0.77 0.90 0.42 0.66 0.49 0.74
Gaussian Surfels (Dai et al., 2024) 0.66 0.93 0.54 0.41 1.06 1.14 0.85 1.29 1.53 0.79 0.82 1.58 0.45 0.66 0.53 0.88
PGSR (Chen et al., 2024a) 0.36 0.57 0.38 0.33 0.78 0.58 0.50 1.08 0.63 0.59 0.46 0.54 0.30 0.38 0.34 0.52
TriangleSplatting (Held et al., 2025a) 0.98 1.07 1.07 0.51 1.67 1.44 1.17 1.32 1.75 0.98 0.96 1.11 0.56 0.93 0.72 1.06
RTS 0.42 0.61 0.74 0.39 0.53 0.86 0.70 0.84 0.72 0.37 0.68 0.87 0.34 0.58 0.44 0.61
@@
Figure 7: Qualitative results on the DTU dataset (Aanæs et al., 2016). RTS estimates high-quality geometry, maintaining thin structures, such as the carrot noses and branches (middle row) and the book edges (bottom row), while effectively modeling smooth surfaces, such as the Buddha (top row).
4.1 Rendering Losses
Following previous work (Kerbl et al., 2023; Huang et al., 2024), we compute the SSIM between the rendered and input images, $\mathcal{L}_{ssim}$, and apply a normal consistency loss (Huang et al., 2024), $\mathcal{L}_{norm}$, to help locally align the triangles with the rendered surface. Please see Section S.3 the supplement.
From the unsupervised depth estimation literature (Chang et al., 2022; Godard et al., 2017; Mahjourian et al., 2018), we adopt a smoothness term on the rendered depth map conditioned on the gradient of the input image. This penalizes large gradients in the rendered depth map where we have small gradients in the input image,
[…]
4.2 Scene Loss
[…]
where $V_{a}^{j}$ and $V_{b}^{j}$ are the $j^{th}$ vertex pair of connected edges $a$ and $b$, respectively. $\widehat{n}_{a}$ and $\widehat{n}_{b}$ are the normals of the connecting triangles and $\Omega$ is the set of all triangles that intersect the current camera frustum. Applying this loss to invisible triangles without rendering losses leads to over-smoothing.
The final objective is a weighted summation of all terms:
5 Experiments
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Ground Truth RTS 2DGS (Huang et al., 2024)
Figure 8: Qualitative comparison between RTS and 2DGS (Huang et al., 2024) on; Top - scan024 from the DTU dataset (Aanæs et al., 2016), Bottom - kitchen from the mip-NeRF 360 dataset (Barron et al., 2022). RTS is substantially more precise at estimating the geometry at discontinuities and rendering fine details, shown in the areas marked by the blue rectangles.
5.1 Implementation Details
We implement the majority of our RTS framework in Python using PyTorch (Paszke et al., 2019). For rasterization, we develop custom CUDA kernels for both the forward and backward pass. We run all our experiments on a single NVIDIA RTX A6000. For our loss weights, we choose $\omega=\begin{bmatrix}1.0&0.05&50.0&1000.0\end{bmatrix}$ empirically.
Following state-of-the-art Multi-View Stereo methods (Yao et al., 2018; Yang et al., 2022; Mi et al., 2022), we directly generate a 3D point cloud for geometric evaluation from the rendered depth maps without performing any TSDF fusion. To generate each point cloud, we use simple heuristic filtering on each depth map, similar to the post-processing presented in GBiNet (Mi et al., 2022). For each depth map, we measure the reprojection error of the depth values at every pixel using neighboring views and filter pixels based on this error. All depth estimates with a low reprojection error are back-projected to 3D points, forming the combined point cloud.
Table 2: Novel View Synthesis on all scenes from the mip-NeRF 360 dataset (Barron et al., 2022).
Method Outdoor Indoor
[…]
NeRF (Mildenhall et al., 2020) 21.46 0.458 0.515 26.84 0.790 0.370
Deep Blending (Hedman et al., 2018) 21.54 0.524 0.364 26.40 0.844 0.261
Instant NGP (Müller et al., 2022) 22.90 0.566 0.371 29.15 0.880 0.216
MipNeRF360 (Barron et al., 2022) 24.47 0.691 0.283 31.72 0.917 0.180
SuGaR (Guédon & Lepetit, 2024) 22.93 0.629 0.356 29.43 0.906 0.225
3DGS (Kerbl et al., 2023) 24.64 0.731 0.234 30.41 0.920 0.189
2DGS (Huang et al., 2024) 24.34 0.717 0.246 30.40 0.916 0.195
GOF (Yu et al., 2024) 24.82 0.750 0.202 30.79 0.924 0.184
PGSR (Chen et al., 2024a) 24.76 0.752 0.203 30.36 0.934 0.147
3DCS (Held et al., 2025b) 24.07 0.700 0.238 31.33 0.927 0.166
TriangleSplatting (Held et al., 2025a) 24.27 0.722 0.217 30.80 0.928 0.160
RTS 21.41 0.657 0.349 30.28 0.921 0.130
5.2 Evaluation
We test our framework on the DTU dataset (Aanæs et al., 2016), an indoor dataset that contains images of 124 scenes taken from a camera mounted on an industrial robot arm. All scenes share the same camera trajectories, with ground-truth point clouds captured via structured light.
We evaluate our new approach on the DTU dataset and record the Chamfer distance in Table 1. We show competitive results alongside the leading state-of-the-art planar GS methods. Across the test set, RTS is the most geometrically accurate in several scenes and second most overall. While PGSR demonstrates impressive reconstruction results, the algorithm utilizes a full suite of multi-view objective functions that significantly improve the geometric reconstruction quality. We provide qualitative results on the DTU dataset in Fig. 7, showing visualizations of the rendered images, depth maps, normal maps, and final point clouds for three scenes. In Fig. 8 (top), we provide a comparison of depth map renderings between RTS and 2DGS (Huang et al., 2024). RTS is able to reconstruct fine details on the surfaces of objects that are typically blurred with Gaussian representations.
Table 3: Ablation study on the contribution of the 3D loss terms using the entire DTU evaluation set. Here, $\lambda_{c}$ is the loss weight for the connectivity term, $L_{conn}$, and $\lambda_{s}$ is the loss weight for the depth smoothness term $L_{smooth}$. The best results are in boldface and the worst are underlined.
@@ 5.2 Evaluation
Method Acc.($m$) $\downarrow$ Comp.($m$) $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ Primitives (K) $\downarrow$
w/o $\mathcal{L}_{smooth}$ & $\mathcal{L}_{conn}$ 0.66 0.68 0.912 30.58 0.227 218
w/o $\mathcal{L}_{smooth}$ 0.65 0.71 0.921 32.13 0.212 244
w/o $\mathcal{L}_{conn}$ 0.67 0.69 0.908 30.15 0.231 224
full ($\lambda_{c}=10.0$, $\lambda_{s}=0.8$) 0.64 0.70 0.918 31.68 0.213 249
full ($\lambda_{c}=300.0$, $\lambda_{s}=20.0$) 0.61 0.63 0.910 30.87 0.232 244
full ($\lambda_{c}=1000.0$, $\lambda_{s}=50.0$) 0.59 0.62 0.909 30.57 0.232 297
We show additional results on the mip-NeRF 360 dataset (Barron et al., 2022). Following the protocol specified by Barron et al. (2022), we separate the images in each scene, taking every eighth image as a test image and training on the remaining. As standard evaluation, we report PSNR, SSIM, and LPIPS (Zhang et al., 2018) metrics. We show results on all indoor and outdoor scenes in Table 2, as well as per-scene results in Table S.1. For indoor scenes, RTS shows results on par with the state-of-the-art methods, having the leading LPIPS score, the fourth overall SSIM score, and a highly competitive PSNR score among all listed methods. On two outdoor scenes (treehill and flowers), RTS is limited in reconstructing extremely distant background foliage, impacting the overall metrics (please see Fig. S.3 for qualitative results on these scenes). In Fig. 8 (bottom), we provide a comparison of novel view synthesis between RTS and 2DGS. RTS is able to render fine textures in low visibility regions in scenes as opposed to blurring with Gaussian representations.
5.3 Ablations
We show an ablation evaluating the contributions of proposed loss terms in Table 3. Removing the soft connectivity leads to a decrease in overall accuracy of the output models, while removing the depth supervision negatively affects the completeness. The two supervision signals complement each other, and we show that increasing the loss weights allows for tuning the framework for either better novel view synthesis or geometry. Additionally, the results corresponding to the changes in magnitude of the loss weights demonstrates the stability of RTS to changes in hyper-parameters.
To evaluate geometry, all previous planar GS algorithms utilize a GT foreground-background segmentation mask when generating the final models. Using this mask to generate the final models for evaluation removes the effects of floaters and inaccurate estimation near the surfaces being evaluated. To portray a more grounded evaluation of geometry, we provide an ablation study in Table S.2 in which we compute Chamfer distance of models without the use of any GT masks.